OpenStudy (anonymous):

The scatter plot below shows the relationship between the length (x), in meters, and mass f(x), in kg, of a marine animal: ordered pairs 1, 1 and 2, 11 and 3, 21 and 4, 31 and 5, 41 and 6, 52 and 7, 65 and 8, 71 and 9, 82 and 10, 94 Which of the following functions is best represented by the scatter plot?

OpenStudy (amistre64):

and our functions are? if you can post a screenshot it may prove useful

OpenStudy (amistre64):

also, define for us what is meant by "best".

OpenStudy (anonymous):

f(x) = 9 − 10.3x f(x) = −10.3 + 9x f(x) = −9 + 10.3x f(x) = 10.3 − 9x

OpenStudy (amistre64):

ok, is the slope going to be positive or negative?

OpenStudy (anonymous):

positive

OpenStudy (amistre64):

then we can rule out 2 options at the start

OpenStudy (anonymous):

wait no the first & last one

OpenStudy (amistre64):

we want to keep b and c

OpenStudy (amistre64):

so, how do we define a reasonable slope? any ideas?

OpenStudy (amistre64):

consider that we move from 1 to 10, i change of 9 we have 2 slopes to work with in our options y/9 = 9 and y/9 = 10.3 it seem to me that y goes from 0 to about 95, which one gives us a y value closest to this?

OpenStudy (anonymous):

y/9=10.3 ?

OpenStudy (amistre64):

i agree

OpenStudy (amistre64):

a slope of 9 gets us to 81, too short, so the 10.3 has to be better

OpenStudy (amistre64):

another sort of test, let x = 10 -10.3 + 9(10) < 90 -9 + 10.3(10) > 90

OpenStudy (anonymous):

-10.3 + 9(10) < 90

OpenStudy (anonymous):

?

OpenStudy (amistre64):

what is 90 - 10 ?

OpenStudy (amistre64):

notice our point at x=10 is about 95, or greater than 90

OpenStudy (anonymous):

f(x) = −10.3 + 9x ?

OpenStudy (amistre64):

if thats your pick for an answer, im going to disagree.

OpenStudy (amistre64):

we want about 95 when x=10 what is -10.3 + 9(10) ?

OpenStudy (anonymous):

its -10.3+9x

OpenStudy (amistre64):

you havent answered my question

OpenStudy (amistre64):

or, are you saying that is what some answer key gives you? you arent being very clear to me.

OpenStudy (anonymous):

im saying that one of the answer choices i thought it was .... sorry

OpenStudy (amistre64):

thats what i had thought you meant but wasnt sure. -10.3 + 9(10) = 79.7 

OpenStudy (amistre64):

do you see that we are not very close to where we need to be?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

-9 + 10.3(10) = 103-9 = 94

OpenStudy (amistre64):

so out of the 2 options f(x) = -9 + 10.3x seems to work better for us

OpenStudy (anonymous):

yea it does

OpenStudy (amistre64):

good luck :)